# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

### v2

@@ Appendix C Derivation of Equation (64)
First, recall that a Clifford unitary is one that preserves the $n$-fold Pauli group $\mathcal{P}_{n}$ when acting on its elements by conjugation [104] (see also [39, Section 3.5]); i.e., the set $C_{n}$ of all Clifford unitaries acting on $n$ qubits is defined as follows:
@@ Appendix D Derivation of Equation (77)
The equalities follow for reasons similar to those given for (71)–(75), with the exception that (256) follows from strong duality of semi-definite programming, under the assumption that there exists at least one state $\rho$ satisfying the constraints in (69). Observe that an optimal $\mu_{0}$ in (257) is equal to the minimum eigenvalue of $H-\mu\cdot Q$. Now suppose that $\rho^{*}$ is optimal for (252) and $\mu^{*}$ is optimal for (257). Then it follows that
@@ G.4 Two-to-four-qubit quantum error-detecting code
The two-to-four qubit quantum error-detecting code encodes two logical qubits into four physical qubits and can detect an arbitrary single-qubit error [61, 62]. Its stabilizer generators are as follows [56, Section 8.1]:
@@
and logical operators for it are as follows [56, Eqs. (8.1)–(8.2)]:

### v3

@@ Appendix C Derivation of Equation (64)
First, recall that a Clifford unitary is one that preserves the $n$-fold Pauli group $\mathcal{P}_{n}$ when acting on its elements by conjugation [119] (see also [47, Section 3.5]); i.e., the set $C_{n}$ of all Clifford unitaries acting on $n$ qubits is defined as follows:
@@ Appendix D Derivation of Equation (77)
The equalities follow for reasons similar to those given for (71)–(75), with the exception that (256) follows from strong duality of semidefinite programming, under the assumption that there exists at least one state $\rho$ satisfying the constraints in (69). Observe that an optimal $\mu_{0}$ in (257) is equal to the minimum eigenvalue of $H-\mu\cdot Q$. Now suppose that $\rho^{*}$ is optimal for (252) and $\mu^{*}$ is optimal for (257). Then it follows that
@@ G.4 Two-to-four-qubit quantum error-detecting code
The two-to-four qubit quantum error-detecting code encodes two logical qubits into four physical qubits and can detect an arbitrary single-qubit error [72, 73]. Its stabilizer generators are as follows [66, Section 8.1]:
@@
and logical operators for it are as follows [66, Eqs. (8.1)–(8.2)]: